OpenStudy (anonymous):

Solve Each System Using Substitution. y-2x=3 3x-2y=5

OpenStudy (anonymous):

Solve for y in the first equation and plug it into the second equation

OpenStudy (anonymous):

but how do i get y= for the first equation givin

OpenStudy (anonymous):

rearrange the equation to get y= ...

OpenStudy (anonymous):

what are the steps to rearrange it

OpenStudy (anonymous):

For the first equation, just plus 2x to both sides to get: y-2x+2x=3+2x The 2x's on the left hand side cancel out as (-2x)+2x=0 Therefore you get y=3+2x.

OpenStudy (anonymous):

Then you just replace the y's in the second equation with 3+2x ...since y=3+2x. And solve.